Automated Recognition of Handprinted Alphanumeric Characters

Sergio Alvarez  Diego Noreña  Edgar Ramos

Department of Electronics Engineering Universidad Javeriana, Bogotá, Colombia

Abstract

We describe the design and implementation of a system for automated recognition of handprinted alphanumeric characters which we have recently developed at the Universidad Javeriana. The system combines structural and statistical elements for character representations, and a Markov model for character transitions in Spanish words. We have constructed a microprocessor–based hardware system to allow character input via an optical pen, and we have built a database of second order transition probabilities in Spanish. In the present paper we also give an overview of various theoretical issues in automated pattern recognition, including discussions of representation, feature extraction, dimensionality reduction, and classification.

Sistemas, Asociación Colombiana de Ingenieros de Sistemas, No. 23, Apr.–Jun. 1985